# Gravitational wave probes of particle dark matter: a review

### v3

@@ V.1 Annihilation signal
When Eq. 94 is no longer satisfied, the cloud stops growing. Since ultralight bosons are their own anti-particles, they will start to annihilate, producing GW radiation at a frequency fixed by the boson mass and slightly redshifted by the BH mass (in the non-relativistic limit $\alpha\ll 1)$, which predominately applies to the equations shown in this work). In the scalar boson case Abbott et al. (2022d):
$f_{\text{gw}}\simeq 483\,{\rm Hz}\left(\frac{m_{\text{b}}}{10^{-12}~{}\text{eV}}\right)\times\left[1-7\times 10^{-4}\left(\frac{M_{\text{BH}}}{10M_{\odot}}\frac{m_{\text{b}}}{10^{-12}~{}\text{eV}}\right)^{2}\right].$ (98)
@@
$\tau_{\rm GW}^{\rm(s)}\approx 2\times 6.5\times 10^{4}\,{\rm yr}\left(\frac{M_{\rm BH}}{10\,M_{\odot}}\right)\left(\frac{0.1}{\alpha}\right)^{15}\frac{1}{\chi_{i}}.$ (99)
$\tau_{\rm GW}^{\rm(v)}\approx 2\times 1\,{\rm day}\left(\frac{M_{\rm BH}}{10\,M_{\odot}}\right)\left(\frac{0.1}{\alpha}\right)^{11}\frac{1}{\chi_{i}}.$ (100)
where we have explicitly noted a factor of 2 that is missing from the original reference Isi et al. (2019), as pointed out in Siemonsen et al. (2023b). There will also be a classical depletion of the cloud over time, as the number of bosons starts to decrease, resulting in the following drifts in frequency over time Isi et al. (2019):
@@
We note that relativistic computations of the frequency drift have also been performed Siemonsen et al. (2023b); May et al. (2025). In the non-relativistic limit $\alpha\ll 1$, the signals will have the corresponding amplitudes of Isi et al. (2019):
@@
$h_{0}^{\rm(v)}\approx 4\times 10^{-24}\left(\frac{M_{\rm BH}}{10M_{\odot}}\right)\left(\frac{\alpha}{0.1}\right)^{5}\left(\frac{\rm Mpc}{r}\right)\left(\frac{\chi-\chi_{f}}{0.1}\right)$ (104)
@@
$\chi_{f}=\frac{4\alpha_{f}m}{4\alpha^{2}_{f}+m^{2}}\,$ (105)
These equations indicate interesting properties for scalar and vector boson cloud annihilation signals. For bosons whose Compton wavelengths are optimally matched to the size of the BH, scalar bosons will annihilate over timescales that greatly exceed the observation time of ground- or space-based GW interferometers, and their frequencies will hardly drift over time. Thus, this signal can be thought of as a quasi-sinusoidal and persistent. On the other hand (and with the same caveat of optimal matching), the GW signal from annihilating vector boson clouds lasts significantly shorter than the scalar one, and thus can be thought of more as a “transient” CW, lasting for durations for $\mathcal{O}($hours-days), i.e. much longer than mergers of black holes routinely detected now by LIGO, Virgo and KAGRA, but much shorter than scalar boson cloud annihilation signals. Furthermore, vector signals emit much more GW power per unit time than scalar ones, but for scalar signals, analysis methods are able to integrate over the entire data collection period of LIGO, Virgo and KAGRA to improve the signal-to-noise ratio. If, in the vector case, the size of the BH does not match the Compton wavelength of the boson, the signals could resemble CWs.
@@ V.2 Transition signal
[…]
$f_{\rm gw}\sim 15\text{ Hz}\left(\frac{m_{b}}{10^{-11}\text{ eV}/c^{2}}\right)$ (106)
@@
and a frequency drift over time that relates to the duration of the signal $T$ and the axion decay constant $f_{a}$ Arvanitaki et al. (2015b):
$\!\!\!\!\frac{df}{dt}\!\!\simeq 10^{-11}\!\frac{\,\mathrm{Hz}\,}{\mathrm{s}}\!\left(\!\frac{f_{\text{gw}}}{90\,\mathrm{Hz}\,}\!\right)\!\!\left(\!\frac{M_{\rm BH}}{10M_{\odot}}\!\right)\!\!\left(\!\frac{10^{17}\!\,\mathrm{GeV}}{f_{a}}\!\right)^{\!2}\!\!\left(\!\frac{5\mathrm{yr}}{T_{\text{obs}}}\!\right)^{\!2}\!\!$ (107)
@@
Unfortunately, the expected event rate of transitions in ground-based detectors is not expected to exceed one even in $\sim 3.5$ years of observing time in advanced LIGO (Fig. 5 in Arvanitaki et al. (2015b)). But, if self-interactions occur, transitions could happen and be detectable Collaviti et al. (2025).
@@
In the vector boson case, transitions could also occur, and the overtone modes could actually grow faster than the fundamental modes at small angular numbers and for fine-structure constants and BH spin values Siemonsen and East (2020). The higher overtones could saturate much quicker than lower ones, implying that transitions to lower states could occur, thus allowing BHs to become unstable to increasing energies at lower modes. Different energy states of the boson cloud could thus be populated at the same time, enriching the kind of physics searchable via transitions.
V.3 Bosenova signal
If the self-interactions between bosons become stronger than the gravitational binding energy of the cloud, the cloud starts to deform and can no longer be described analogously to the Hydrogen atom Fukuda and Nakayama (2020); Collaviti et al. (2025); Omiya et al. (2024); Jones et al. (2025). Instead, it can collapse and explode in what is known as a “bosenova”. This phenomenon would occur if the potential energy of the cloud is approximately equal to the self-interaction energy. Bosenovas have been observed on earth in condensed matter laboratories, in which repulsive forces were switched to attractive ones in a material that formed a Bose-Einstein condensate via a tuneable external magnetic field Donley et al. (2001). There are a lot of uncertainties around whether the bosenova would occur, and it seems that including self-interactions in the cloud limits bosenovas to a corner of the allowed parameter BH-boson mass parameter space Baryakhtar et al. (2021).
@@ V.3 Bosenova signal
Though GWs from boson clouds described above may seem independent, they could in principle be happening at the same time. In particular, if ultralight scalar or vector bosons have some self-interactions, the GW signal will change. If these clouds are composed of string axions (as opposed to the QCD axion), a non-linear self-interaction may cause a bosenova if, during superradiance, the value of the scalar field approaches the axion decay constant $f_{a}$, which, for string axions, could be below the grand unification theory (GUT) scale Arvanitaki et al. (2010); Yoshino and Kodama (2012, 2015a, 2015b). This bosenova would halt the superradiance process, as about 5% of the boson cloud mass would fall back into the BH, but superradiance would immediately start again until the next bosenova is triggered. Such periodic bosenovas could sustain superradiance for much longer durations than those in the conventional scenario, permitting annihilating clouds even around extremely old BHs. Furthermore, we note that boson annilation could alter the GW of an inspiraling binary, potentially visible in space-based GW detectors Xie and Huang (2024), in a similar way that dynamical friction does.
In more general models of axions than in the axiverse, self-interactions could suppress annihilation signals as they get stronger, and transitions could occur for moderate interaction strengths. In particular, GW power sourced by annihilation would be reduced due to the smaller cloud size that can be sustained because of self-interactions, while scalar radiation, a new energy loss mechanism, consumes a portion of the energy that would source GW emission. But, multiple levels could be populated with self-interactions, allowing transitions to occur more than in the non-self-interaction scenarios Baryakhtar et al. (2021).
V.4 Clouds in binary black hole systems
@@
We know that BHs form binaries, and if ultralight dark matter exists in the form of ultralight bosons, it should imprint some signature in the GWs arising from inspiraling binary black holes Baumann et al. (2019a). The GW signal will be modified due to tidal disruptions induced by the companion, the multiple moments of the cloud Baumann et al. (2019a, b, 2020) or through a dynamical friction drag forced induced by the ultralight dark matter. These kinds of signals could be visible in future space-based detectors, since the deviations primarily occur in the early inspiral stages of the system, i.e. at frequencies much lower than currently accessible by ground-based GW interferometers. In both cases, observations of the inspiral of binary systems, and in particular extreme mass ratio inspiral (EMRI) systems Zhang and Yang (2019); Brito and Shah (2023), could indicate the presence of boson clouds around one or both of the objects if there is some dephasing with respect to the vacuum . Additionally, once the binary reaches a certain stage of its evolution, the impact of both the spin-induced quadrupole moment and tidal forces will lessen, since the clouds will become disrupted by their companions. If we can observe variations in both the spin-induced quadrupole moment, parameterized as $\kappa$, and the tidal deformability, parameterized as $\Lambda$, over the course of the lifetime of the binary, it would indicate the presence of a (disrupted) boson cloud Baumann et al. (2019a). More recently, it has been show that the boson cloud could be “ionized” as the orbital separation approaches the size of the cloud. Self-interactions could also play a role in the evolution of a binary, in which two types of clouds can actually coexist around the rotating BH through mode coupling Boudon et al. (2024); Aurrekoetxea et al. (2024a); Takahashi et al. (2024). We will now describe different ways in which boson clouds could affect GWs from binary black hole systems.
V.4.1 Spin-induced quadrupole moment
@@
The spin-induced quadrupole moment is often parameterized in terms of a dimensionless quantity $\kappa=\frac{Q}{J^{2}/M_{\rm BH}}$, which has been normalized by the object’s angular moment $J$ and mass BH. This parameter enters into the waveform at the second post-Newtonian order. Thus, the spin-induced quadrupole moment, regardless of its source, has a significant impact on the GW signal from compact objects inspiraling towards one another.
@@ V.4.2 Tidal disruptions
Because the boson cloud is much less compact than the BH, we expect that tidal disruptions from the companion could occur when the two objects in the binary black hole are close enough to each other Ng et al. (2020). To characterize the effects of tides in a binary system, the tidal deformability parameter $\Lambda$ is used, which depends on the mass and radius of the compact objects. For BHs, this quantity is 0, while for neutron stars, it can take on values of $\mathcal{O}(100-1000)$ Chatziioannou (2020). For boson clouds, $\Lambda\sim 10^{7}$, which is a remarkable departure from that expected from conventional astrophysical binaries Baumann et al. (2019b). In Payne et al. (2022), the authors considered the impact that superradiance would have on hierarchical BH mergers, and showed that the reduction of spin in BHs implies smaller recoil velocities of binary black hole mergers. Smaller recoil velocities imply that remnants could remain inside clusters, leading to higher chances of hierarchical formation and merger of binary black holes. Such disruptive effects of bosons on not just individual BHs but also on the BH population, imply that boson clouds, if they exist, would significantly impact the formation and evolution histories of BHs.
V.4.3 Binary-induced transitions
Over the course of the inspiral, the presence of the boson cloud will induce transitions between different energy levels of the boson clouds that would be visible in space-based GW detectors if (1) the orbital frequency matches the hyperfine splitting between two energy levels (hyperfine resonance), and (2) the growing mode mixes with decaying modes (Bohr resonance). Both kinds of transitions would also be smoking-gun signatures of the presence of boson clouds around rotating BHs in a binary Baumann et al. (2019a). More recently, by considering the slow evolution of the orbital frequency through resonances of the boson cloud system, that is, when the orbital frequency matches the energy spacing between adjacent levels, it can be shown that Landau-Zener transitions occur and significantly affect the inspiral GW signal Baumann et al. (2020).
V.4.4 Ionization

### v4

@@ V.1 Annihilation signal
Ultralight bosons are their own anti-particles, which means that they continuously annihilate (even as the cloud is growing), producing GW radiation at a frequency fixed by the boson mass and slightly redshifted by the BH mass (in the non-relativistic limit $\alpha\ll 1$, which predominately applies to the equations shown in this review). In the scalar boson case Abbott et al. (2022d):
$f_{\text{gw}}$ $\simeq 483\,{\rm Hz}\left(\frac{m_{\text{b}}}{10^{-12}~\text{eV}}\right)$
$\times\left[1-7\times 10^{-4}\left(\frac{M_{\text{BH}}}{10M_{\odot}}\frac{m_{\text{b}}}{10^{-12}~\text{eV}}\right)^{2}\right].$ (101)
@@
$\tau_{\rm GW}^{\rm(s)}\approx 2\times 6.5\times 10^{4}\,{\rm yr}\left(\frac{M_{\rm BH}}{10\,M_{\odot}}\right)\left(\frac{0.1}{\alpha}\right)^{15}\hskip-2.0pt\frac{1}{\chi_{i}}.$ (102)
$\tau_{\rm GW}^{\rm(v)}\approx 2\times 1\,{\rm day}\left(\frac{M_{\rm BH}}{10\,M_{\odot}}\right)\left(\frac{0.1}{\alpha}\right)^{11}\hskip-2.0pt\frac{1}{\chi_{i}}.$ (103)
where we have explicitly noted a factor of 2 that is missing from the original reference Isi et al. (2019), as pointed out in Siemonsen et al. (2023b).
As the number of bosons starts to decrease, there will also be a classical depletion of the cloud over time. This reduction in bosons causes the cloud to contract, resulting in the following positive drifts in frequency over time Isi et al. (2019) [^11] We note that relativistic computations of the frequency drift have also been performed Siemonsen et al. (2023b); May et al. (2025), which are useful for real searches of GWs arising from boson clouds but not less important for our discussions here.:
@@
In the non-relativistic limit $\alpha\ll 1$, the signals will have the corresponding amplitudes of Isi et al. (2019):
@@
$h_{0}^{\rm(v)}\approx 4\times 10^{-24}\left(\frac{M_{\rm BH}}{10M_{\odot}}\right)\left(\frac{\alpha}{0.1}\right)^{5}\left(\frac{\rm Mpc}{r}\right)\left(\frac{\chi-\chi_{f}}{0.1}\right),$ (107)
@@
$\chi_{f}=\frac{4\alpha_{f}m}{4\alpha^{2}_{f}+m^{2}}\,.$ (108)
These equations indicate interesting properties for scalar and vector boson cloud annihilation signals. For bosons whose Compton wavelengths are optimally matched to the size of the BH, scalar bosons will annihilate over timescales that greatly exceed the observation time of ground- or space-based GW interferometers (Eq. 102), and their frequencies will hardly drift over time. Thus, this signal can be thought of as a quasi-sinusoidal and persistent. On the other hand (and with the same caveat of optimal matching), the GW signal from annihilating vector boson clouds lasts significantly shorter than the scalar one (Eq. 103), and thus can be thought of more as a “transient” CW, lasting for durations for $\mathcal{O}($hours-days), i.e. much longer than mergers of black holes routinely detected now by LIGO, Virgo and KAGRA, but much shorter than scalar boson cloud annihilation signals. Furthermore, vector signals emit much more GW power per unit time than scalar ones, but for scalar signals, analysis methods are able to integrate over the entire data collection period of LIGO, Virgo and KAGRA to improve the SNR, as will be discussed in Section VI.2. However, in the vector case, it is possible to have CW emission for particular combinations of boson mass and BH mass, such as large BHs and small boson masses, where the spin-up of the signal, as described by Eq. 105, becomes comparable to that in the scalar case (Eq. 104), thus making the signal almost monochromatic.
@@ V.2 Transition signal
In the scalar boson case, just as electrons jump between discrete energy levels in atoms, bosons can also transition between discrete energy states, provided these states are populated. The energy states of the boson cloud are quantized, with each state corresponding to specific quantum numbers. Significant transition signals are expected when the cloud reaches a population that allows transitions between particular energy levels, such as from the 5g to the 6g states, where “g” refers to the orbital angular momentum quantum number of the state. These transitions between specific energy levels are expected to generate detectable gravitational wave signals that could last for durations much longer than the typical observation times of ground-based gravitational wave interferometers. The frequency of these signals is given by Arvanitaki et al. (2015a):
$f_{\rm gw}\sim 15\,\text{Hz}\left(\frac{m_{b}}{10^{-11}\,\text{eV}}\right)$ (109)
@@
The frequency of the signal would drift over time according to the duration of the signal and the axion decay constant $f_{a}$ Arvanitaki et al. (2015a):
$\frac{df}{dt}\simeq 10^{-11}\,\frac{\text{Hz}}{\text{s}}\left(\frac{f_{\rm gw}}{90\,\text{Hz}}\right)\left(\frac{M_{\rm BH}}{10\,M_{\odot}}\right)\left(\frac{10^{17}\,\text{GeV}}{f_{a}}\right)^{2}\left(\frac{5\,\text{yr}}{T_{\rm obs}}\right)^{2}$ (110)
We note here that the axion decay constant, arising from self-interactions, affects the spin-up of the signal. In the previous discussions surrounding Eqs. 104 and 105, self-interactions were assumed to be zero.
@@
While the expected event rate of these transitions in ground-based detectors is unlikely to exceed one event in around 3.5 years of observation with advanced LIGO Arvanitaki et al. (2015a), the presence of self-interactions could potentially enhance the transition rates, making these signals detectable Collaviti et al. (2025).
@@
In the vector boson case, transitions can also occur, and interestingly, overtone modes may grow faster than the fundamental modes, particularly for small angular numbers, fine-structure constants, and black hole spin values Siemonsen and East (2020). These higher overtones could saturate more quickly than the fundamental modes, allowing transitions to lower states and increasing the overall energy of the cloud, potentially leading to detectable gravitational wave emission.
V.3 Impact of boson self-interactions on superradiance
While the formation of ultralight boson clouds around rotating BHs is already expected to result in observable GWs, the strength of the self-interactions within the boson cloud can play a crucial role in determining the nature and detectability of these signals. Specifically, the self-interactions between bosons in the cloud can influence key dynamical processes, such as:
  1. 1.
    Saturation of Superradiance: Self-interactions modify the growth timescale, potentially leading to a faster saturation of the superradiance process. This could result in stronger GW emission once the cloud reaches its maximum energy state. The timescale for superradiance saturation is sensitive to the coupling strength between bosons, and self-interactions may allow the cloud to reach a critical size faster, enhancing the CW signal emitted during this phase Yoshino and Kodama (2012, 2015a, 2015b).
  2. 2.
    Merger Dynamics: The presence of self-interactions can affect how quickly the boson cloud dissipates or decays, influencing the coalescence of the black hole system. This decay modifies the timescale of the inspiral phase and the subsequent merger, altering the frequency evolution of the emitted GWs. A faster decay due to self-interactions could lead to a sharper rise in the frequency of the GW signal Boudon et al. (2024); Aurrekoetxea et al. (2024); Takahashi et al. (2024).
  3. 3.
    Enhanced Annihilation of Boson Clouds in Isolated Black Holes: In a non-interacting system, the annihilation rate of bosons is limited by the density and distribution of particles in the cloud. Self-interactions, however, can increase the density of the cloud by accelerating the process of boson accumulation and scattering. As a result, the rate of boson annihilation can be significantly enhanced. This increased annihilation rate leads to stronger and potentially more detectable GW signals from the cloud Baryakhtar et al. (2021).
  4. 4.
    Enhanced Energy-Level Transitions: The energy levels of bosons in the cloud are quantized, similar to the energy levels of electrons in an atom. Self-interactions can facilitate the transition of bosons between higher energy states more rapidly, increasing the amplitude of the emitted gravitational waves. This could make energy-level transitions a more significant source of detectable GW signals, especially if the cloud is in a high-energy state Baryakhtar et al. (2021).
These effects become particularly relevant when considering the potential detection of boson clouds in future GW observatories, such as Cosmic Explorer and ET Collaviti et al. (2025). Recent searches for these signals have not accounted for these interactions, and their inclusion may shift current detection prospects by enhancing the signal strength or changing the emission characteristics. In this section, we will explore how self-interactions influence the dynamics of boson clouds and their implications for gravitational wave astronomy.”
V.4 Bosenova signal
If the self-interactions between bosons become stronger than the gravitational binding energy of the cloud, the cloud begins to destabilize and cannot be described analogously to the hydrogen atom Fukuda and Nakayama (2020); Collaviti et al. (2025); Omiya et al. (2024); Jones et al. (2025). The cloud may undergo a significant reconfiguration, leading to a dramatic release of energy, a process referred to as a “bosenova.” This occurs when the potential energy of the boson cloud becomes comparable to or exceeds the self-interaction energy, leading to the rapid rearrangement of the bosonic matter.
The phenomenon of a bosenova has been observed in condensed matter systems, in which a Bose-Einstein condensate undergoes a similar process due to a change in interactions induced by an external magnetic field Donley et al. (2001). However, the occurrence of a bosenova in the context of ultralight bosons around black holes remains highly uncertain, with studies suggesting that self-interactions may limit bosenovas to a restricted region of the allowed parameter space, particularly for certain combinations of black hole and boson masses Baryakhtar et al. (2021).
@@ V.4 Bosenova signal
Though GWs from boson clouds described above may seem independent, they could in principle be happening at the same time. In particular, if ultralight scalar or vector bosons have some self-interactions, and If these clouds are composed of string axions (as opposed to the QCD axion), a non-linear self-interaction may cause a bosenova if, during superradiance, the value of the scalar field approaches the axion decay constant $f_{a}$, which could be below the grand unification theory (GUT) scale Arvanitaki et al. (2010); Yoshino and Kodama (2012, 2015a, 2015b). This bosenova would halt the superradiance process, as about 5% of the boson cloud mass would fall back into the BH, but superradiance would immediately start again until the next bosenova is triggered. Such periodic bosenovas could sustain superradiance for much longer durations than those in the conventional scenario, e.g. in Arvanitaki and Dubovsky (2011), permitting annihilating clouds even around extremely old BHs, and even affect GWs from binary black hole systems Xie and Huang (2024) .
V.5 Clouds in binary black hole systems
@@
If boson clouds form around binary black hole systems, they should imprint some signature in the GWs arising from their inspiral Baumann et al. (2019a). In particular, the GW signal will be modified due to tidal disruptions induced by the companion, the multiple moments of the cloud Baumann et al. (2019a, b, 2020) or through a dynamical friction drag force induced by the boson clouds. These kinds of signals could be visible in future space-based detectors, since the deviations primarily occur in the early inspiral stages of the system, i.e. at frequencies much lower than currently accessible by ground-based GW interferometers Guo et al. (2024a). In both cases, observations of the inspiral of binary systems, and in particular EMRI systems Zhang and Yang (2019); Brito and Shah (2023), could indicate the presence of boson clouds around one or both of the objects if there is some dephasing with respect to the waveform in vacuum. Additionally, once the binary reaches a certain stage of its evolution, the impact of both the spin-induced quadrupole moment and tidal forces will lessen, since the clouds will become disrupted by their companions. If we can observe variations in both the spin-induced quadrupole moment and the tidal deformability over the course of the lifetime of the binary, it would indicate the presence of a (disrupted) boson cloud Baumann et al. (2019a).
More recently, it has been shown that as the orbital separation in a binary black hole system decreases and becomes comparable to the size of the boson cloud, the boson cloud could be “ionized” Takahashi et al. (2022); Baumann et al. (2022a, b). This ionization occurs when the energy from the binary black hole system becomes large enough to unbind the cloud from its host BH. As a result, the inspiral of the binary becomes significantly influenced by the dynamics of the boson cloud, rather than merely being perturbed by it.
We will now describe four different ways in which boson clouds could affect GWs from binary black hole systems.
V.5.1 Spin-induced quadrupole moment
@@
The spin-induced quadrupole moment is often parameterized in terms of a dimensionless quantity $\kappa=\frac{Q}{J^{2}/M_{\rm BH}}$, where $Q$ is the quadrupole moment of the BH, that has been normalized by the angular moment $J$ and mass of the BH. This parameter enters into the waveform at the second post-Newtonian (2PN) order, i.e. as corrections of order $(v/c)^{4}$ relative to the leading quadrupole radiation. Thus, the spin-induced quadrupole moment, regardless of its source, has a significant impact on the GW signal from compact objects inspiraling towards one another.
@@ V.5.2 Tidal disruptions
Because the boson cloud is much less compact than the BH, we expect that tidal disruptions from the companion could occur when the two objects in the binary black hole are close enough to each other Ng et al. (2020). To characterize the effects of tides in a binary system, the tidal deformability parameter $\Lambda$ is used, which depends on the mass and radius of the compact objects. For BHs in vacuum, this quantity is 0, while for neutron stars, it can take on values of $\mathcal{O}(100-1000)$ Chatziioannou (2020). For boson clouds, $\Lambda\sim 10^{7}$, which is a remarkable departure from that expected from conventional astrophysical binaries Baumann et al. (2019b). In Payne et al. (2022), the authors considered the impact that superradiance would have on hierarchical BH mergers, and showed that the reduction of spin in BHs implies smaller recoil velocities of binary black hole mergers. Smaller recoil velocities imply that remnants could remain inside clusters, leading to higher chances of hierarchical formation and merger of binary black holes. Such disruptive effects of bosons on not just individual BHs but also on the BH population, imply that boson clouds would significantly impact the formation and evolution histories of BHs.
V.5.3 Binary-induced transitions
During the inspiral of a binary black hole system that contains boson clouds, the interaction between the cloud and the BHs can cause transitions between different energy levels in the clouds. These transitions can potentially be detected by space-based GW interferometers if any of the following mechanisms is present:
  1. 1.
    Hyperfine Resonance: The orbital frequency of the binary system matches the energy spacing between two adjacent energy levels in the boson cloud. This condition is analogous to a resonance in atomic physics, where the energy difference between two states is exactly matched by an external driving frequency. In this case, the system stays in equilibrium, and energy is transferred between energy levels that are in resonance Baumann et al. (2019a).
  2. 2.
    Bohr Resonance: In addition to hyperfine resonance, the growing mode (the mode that is being excited by the inspiral) mixes with decaying modes of the cloud. This interaction allows energy to be transferred between different states, influencing the GW signal Baumann et al. (2019a).
  3. 3.
    Landau-Zener Transitions: When the orbital frequency of the inspiral matches the energy spacing between adjacent energy levels, the system undergoes a non-adiabatic shift in energy states. This transition process, characterized by the Landau-Zener mechanism, significantly alters the frequency and amplitude evolution of the GW signal. These non-adiabatic transitions can provide observable changes in the signal, offering additional insights into the properties of the boson cloud Baumann et al. (2020).
These effects, all of which could be present simultaneously, would provide “smoking-gun” signatures of the presence of boson clouds around rotating black holes in a binary system Baumann et al. (2019a).
V.5.4 Ionization